OpenStudy (anonymous):

prove that : cos64 + sin 64 * tan32 = 1

OpenStudy (hoblos):

cos64 = cos[2(32)] = cos^2 (32) - sin^2 (32) sin 64 = sin[2(32)] = 2sin32cos32 tan32 = sin32 / cos32 cos64 + sin 64 * tan32 = cos^2 (32) - sin^2 (32) + 2sin32cos32* sin32 / cos32 = cos^2 (32) - sin^2 (32) + 2sin^2 (32) =cos^2 (32) + sin^2 (32) =1

OpenStudy (anonymous):

thank you so much dude :) i really appreciate it

OpenStudy (hoblos):

my pleasure XD

OpenStudy (anonymous):

\[\tan ^2x =( \cos (180-x)\sin(x-90)-1)\[\div\]tan^2(180+x)\sin(90+x)\cos(-x)\]